OpenStudy (anonymous):

Calculate the boiling point of a solution prepared by dissolving 70.0 g of naphthalene, C10H8 (a nonvolatile nonelectrolyte), in 220.0 g of benzene, C6H6. The Kb for benzene = 2.53oC/m. The boiling point of pure benzene is 80.1oC. Ans: 86.4 degrees Celsius I did this so far 1)70G C10H8(1mol c10H8/128gC10H8)= .546mol C10H8 2)(.546mol c10h8)(.220kg benzene)=2.482m 3) (2.482)(2.53 C/m)=6.289 I'm not sure if I am starting this off right, can anyone help me get the correct answer? Please ans ty!

OpenStudy (chmvijay):

ΔT =(Kb×1000×W2)/M2×W1 = 2.53 *1000*70 /128.17 * 220 =177100 /28197.4 =6.28 but ΔT=T-To 6.28 =T - 80.1 T=6.28+80.1 =86.38 °C

OpenStudy (chmvijay):

where W2 - gram of non volatile solute naphthalene W1 - gram of solvent benzene, M2 molar mass of naphthalene ,